OpenStudy (anonymous):

sin 5 theta?using theorem 5

Parth (parthkohli):

\[\sin(3\theta + 2\theta) = \sin(3\theta) \cos(2\theta) + \sin(2\theta)\cos(3\theta)\]

Parth (parthkohli):

\[\cos(2\theta) = 1 - 2\sin^2 \theta\]\[\sin(3\theta) = 3\sin\theta - 4\sin^3\theta\]\[\sin(2\theta) = 2\sin\theta \cos \theta\]\[\cos(3\theta) = \cos\theta (4\cos^2 \theta - 3)\]